OpenStudy (anonymous):

Use Part 1 of the Fundamental Theorem of Calculus to find the derivative of the function.

OpenStudy (anonymous):

u = 1/x

OpenStudy (anonymous):

let \[ f'(t) = \arctan t \]Then \[ \int_7^uf'(t)\;dt = f(u) - f(7) \]

OpenStudy (anonymous):

\[ \frac{d}{du}\left[f(u)-f(7)\right] = f'(u) = \arctan u \]Remember that \(f'(7) =\frac d{du} \arctan 7 = 0\)

OpenStudy (anonymous):

That is the FTC part.

OpenStudy (anonymous):

The other part is simply: \[ \frac{du}{dx} = \frac{d}{dx}u = \frac{d}{dx}\left(\frac 1x\right) \]

OpenStudy (anonymous):

what should I put in these two space

OpenStudy (anonymous):

Actually, I shouldn't have written \(f'(7)\), that is notation failure. It should remain as \(\frac{d}{du}f(7)\).

OpenStudy (anonymous):

If you understand what I said, you know what to do.